OpenStudy (anonymous):

2ndderivative of xcosa+ysina=p,

OpenStudy (anonymous):

I must say any more information

OpenStudy (anonymous):

that's just the given.. eliminating the arbitrary constant is the goal..

OpenStudy (anonymous):

ok ok

OpenStudy (agent47):

ySin(a)=p-xCos(a) y=(1/Sina)(p-xCos(a)) I'm assuming x is the only variable.

OpenStudy (agent47):

y'=(1/Sina)(-Cos(a)) y''=0

OpenStudy (agent47):

or maybe it's p(a)

OpenStudy (agent47):

What is it? y(x), y(a), x(a), p(x), p(y), etc... you choose

OpenStudy (anonymous):

well it could be y-2=xcos(a)+y-2=0

OpenStudy (agent47):

We'll differentiate!

OpenStudy (anonymous):

umm....

OpenStudy (anonymous):

o isee now it is ySin(a)=p-xCos(a) y=(1/Sina)(p-xCos(a)

OpenStudy (anonymous):

y''sina=0